stu169947:

What is the result of factoring out the GCF from the expression (24 + 36)

dude:

To find the GCF you first have to know the factors of each 24 1\(\times\)24 2\(\times\)12 3\(\times\)8 4\(\times\)6 36 1\(\times\)36 2\(\times\)18 3\(\times\)12 4\(\times\)9 6\(\times\)6 Look for the greatest number they share and take it out of the parenthesis and multiply their respective factor in the parenthesis